# A Detection of Circumgalactic Dust at Megaparsec Scales with Maximum Likelihood Estimation

### v1

@@ Abstract
One of the more surprising astrophysical discoveries of the last decade has been the presence of enormous quantities of dust at megaparsec distances from galaxies, which has important implications for galaxy evolution, the circumgalactic and intergalactic medium, and observational cosmology. In this work, we present a novel method for studying these vast halos of circumgalactic dust: a maximum-likelihood estimator for dust-induced extinction of background galaxies. This estimator can accommodate a broad range of archival photometric data and incorporate different dust reddening prescriptions, making it applicable to diverse galaxy types and redshifts. We apply the estimator to the redMaGiC catalog of luminous red galaxies, selected for their tight dispersion in color and well-constrained photometric redshifts, and measure the resulting extinction as a function of projected distance from WISExSuperCOSMOS and redMaGiC foreground galaxies. We detect significant dust-induced extinction profiles extending to at least 1 megaparsec from galactic disks, with noticeable differences between star-forming and quiescent galaxies: star-forming galaxies exhibit a pronounced rise in extinction within the inner 50 kiloparsecs and a steep decline beyond 1 megaparsec, while the quiescent galaxies host little dust in the inner halo but have detectable extinction out to 30 megaparsecs. We test the robustness of our results using star catalogs and inverted foreground and background samples and find no evidence for significant systematic error. Our approach provides a powerful tool for studying the interplay between circumgalactic dust, galaxy evolution, and large-scale structure, with potential applications in a number of astrophysical subfields.
software: Astropy (Astropy Collaboration et al., 2013, 2018), NumPy (C. R. Harris et al., 2020), TreeCorr (M. Jarvis et al., 2004), healpy (A. Zonca et al., 2019)
@@ 1 Introduction
Most of the universe’s atoms are hidden from plain sight: galaxies account for less than 10% of baryons, while the remaining 90% reside in gaseous phases whose low densities make them difficult to observe. These include the circumgalactic medium (CGM), which permeates the dark matter halos hosting galaxies, and the more extended intergalactic medium (IGM) that connects them (M. Fukugita & P. J. E. Peebles, 2004; J. M. Shull et al., 2012). Combined, they are the reservoir for more than 85% of the baryons in the universe (F. Walter et al., 2020). Despite being the focus of intense theoretical and observational research, many open questions remain in the study of CGM and IGM, particularly regarding the origin, distribution, and survival of cool, condensed matter in these environments.
@@
One pressing issue is the existence of huge amounts of dust in the CGM and IGM—comparable to the amount of dust contained within galaxies themselves—which challenges models of the cosmic baryon cycle for several reasons (S. Veilleux et al., 2020; C. Péroux & J. C. Howk, 2020; H.-W. Chen & F. S. Zahedy, 2024). First, dust is predominantly deposited into the interstellar medium (ISM) through supernovae and asymptotic giant branch star winds; second, the CGM and IGM are filled with mega-Kelvin gas, which should rapidly destroy dust grains via sputtering (B. T. Draine & E. E. Salpeter, 1979). While a number of mechanisms for dust deposition and replenishment have been proposed, including stellar radiation-pressure driven outflows, tidal stripping of interacting galaxies, supernova-driven outflows, and cool outflows, the exact coupling between the multi-phase CGM, IGM, and ISM remains unclear (H.-W. Chen & F. S. Zahedy, 2024).
@@
Understanding this connection matters because the abundance and spatial extent of dust in the CGM is a key component of galaxy formation and evolution models. Dust also fundamentally limits the homogeneity of samples in wide-field surveys and introduces significant scatter to cosmological distance measurements, particularly supernovae. According to D. Rubin et al. (2023), dust is the second-largest source of uncertainty in Type Ia supernova cosmological parameter constraints, accounting for 11.5% of the total variance. Current analyses include intergalactic dust reddening in the systematic error budget, but because the magnitude of the effect is uncertain, they do not correct the distance measures themselves.
@@
The existence of circumgalactic dust at megaparsec scales was first reported by Ménard, Kilbinger, & Scranton (2010), hereafter M10, who found a smooth extinction profile extending to 10 Mpc while measuring the magnification of 85,000 background quasars by 20 million galaxies in SDSS images. A similar measurement was carried out by J. E. G. Peek et al. (2015) using a background sample of luminous red galaxies (LRGs), whose small color dispersion enabled precise measurement of reddening by dust halos with a smaller sample size than M10. However, the majority of studies of dust in the extended CGM have been spectroscopic, e.g., G. Zhu et al. (2014), T.-W. Lan & H. Mo (2018), Y. Zu (2021), and Z. Chen et al. (2024). These studies typically use tracers like the equivalent width (EW) of Mg II absorption lines to probe the dynamics and shape of dust halos of galaxies at $z>0.3$, for which the Mg II $\lambda\lambda$ 2796, 2803 lines are redshifted into the visible. With the notable exception of Z. Ruoyi & Y. Haibo (2020), who report the detection of dust-induced reddening about M31 and M33 at projected distances greater than 100 kpc, few measurements have replicated the original circumgalactic dust measurement using extinction of background sources.
@@
These observational gaps highlight the need for a way to assess dust in the extended CGM without relying on atomic absorption lines ill-suited for lower-redshift galaxies, particularly given the evidence that the cosmic dust density changes significantly at $z<1$ (C. Péroux & J. C. Howk, 2020). Drawing inspiration from the “standard crayon” approach of J. E. G. Peek et al. (2015) and from analytical techniques ubiquitous in cosmology, we present a new technique to detect circumgalactic dust halos: a maximum-likelihood estimator for dust-induced extinction with a lower bound on the variance set by the Cramér-Rao bound. Application of the estimator is relatively simple, requiring only a photometric dataset and specification of a theoretical model for dust-induced reddening. Applying this estimator to the redMaGiC galaxy catalog, notable for its tight dispersion in color, and cross-correlating the inferred extinction with $z<0.2$ galaxies from the WISExSuperCOSMOS catalog, we find fresh evidence for circumgalactic dust halos at distances of 1 Mpc $h^{-1}\,$or greater from the galactic disk. We also analyze a sample of $0.450.15<z<$ 0.45 redMaGiC LRGs, and find distinct differences in the radial extinction profiles of star-forming and quiescent galaxies.
@@
The remainder of this paper is organized as follows. We describe our dust extinction estimator in Section 2, provide details on datasets in Section 3, and describe our methods in Section 4. Results are shown in Section 5 and discussed in Section 6. We conclude in Section 7.
2 A maximum likelihood estimator for dust extinction
@@
A screen of dust absorbs and scatters incoming light, dimming and reddening the spectrum. The extinction $A_{\lambda}$ quantifies wavelength-dependent attenuation of light due to dust; reddening is usually parametrized by the selective extinction of red versus blue light in the UBV photometric system: $E(B-V)$. The steepness of $\rm A_{\lambda}$ as a function of wavelength is parametrized by the ratio of selective-to-total extinction in $V$-band: $R_{V}=A_{V}/E(B-V)$. For interstellar Milky Way dust, $R_{V}\sim 3.1$; dense molecular clouds have $R_{V}\sim 5$ (E. L. Fitzpatrick & D. Massa, 2007).
@@
Most photometric studies of the CGM trace dust using some variant of the selective extinction $E(B-V)$; here, we use maximum likelihood estimation (e.g., R. J. Barlow, 1993) to find a parameter that represents the best chance of finding dust-induced extinction based on a change in background galaxy colors. This estimator can be derived by maximizing the log-likelihood
@@
and the covariance of the data $\mathbf{C}$ encapsulates both the magnitude measurement uncertainties and cosmic variance. The model $\mathbf{M}(\rho)$ describes the data $\mathbf{D}$ in terms of a parameter $\rho$ (here, $A_{V}$); the model can be Taylor-expanded to obtain a simple linear approximation for the effect of $\rho$:
@@
In this expansion, $\mathrm{M_{0}}$ are the intrinsic colors of galaxies in the background galaxy sample, without any reddening from dust. The $\rho$ $d\mathrm{M}/d\rho$ term is the response of galaxies to a differential change in extinction; $d\mathrm{M}/d\rho$ is a fiducial extinction curve for dust with a given $R_{V}$. With this assumption, the optimal estimator $\hat{\rho}$ for dust-induced extinction $\rho$ such that $\rho$ maximizes the likelihood in Equation 1 is given as:
@@
With the extinction estimator $\hat{\rho}$ in place, foreground galaxy number counts may be cross-correlated with the dust-induced extinction of background galaxies. Denoting $N(\theta_{1})$ as the galaxy count at a position $\theta_{1}$ on the sky and $A_{V}(\theta_{2})$ as the extinction at position $\theta_{2}$, the cross-correlation may be expressed as the average deviations of $N$ and $A_{V}$ from their mean values at two sky positions $\theta_{1}$ and $\theta_{2}$ separated by $\theta$:
$\xi_{NA_{V}}(\theta)=\left<[N(\theta_{1})-\overline{N}]\,[A_{V}(\theta_{2})-\overline{A_{V}}]\right>$ (6)
@@
The mean values, $\overline{N}$ and $\overline{A_{V}}$, are the average foreground galaxy count and average extinction of background galaxies, respectively, over the entire survey area.
@@
As with any such analysis, the contribution of random correlations to $A_{V}$ must be taken into account. We use the following estimator for the true galaxy-reddening cross-correlation $\xi_{NA_{V}}(\theta)$:
$\xi_{NA_{V}}(\theta)=D_{N}D_{A_{V}}(\theta)-D_{N}R_{A_{V}}(\theta)-R_{N}D_{A_{V}}(\theta)+R_{N}R_{A_{V}}(\theta)$ (7)
@@
$D_{N}D_{A_{V}}$ is the signal from the correlation of the real foreground galaxy and background color catalogs; $D_{N}R_{A_{V}}$ and $R_{N}D_{A_{V}}$ are the correlations of real foreground with random background catalog and random foreground with real background catalogs, respectively; and $R_{N}R_{A_{V}}$ is the correlation of the random foreground and background catalogs.
3 Data
@@
We use the Dark Energy Survey (DES) Y3 redMaGiC catalogs (E. Rozo et al., 2016; S. Pandey et al., 2022) as background samples and the WISExSuperCOSMOS photometric redshift catalog (M. Bilicki et al., 2016), plus a low-redshift subset of redMaGiC galaxies, as foreground samples. In addition, star catalogs selected from the Gaia survey and DES Y3 are used to test for systematic error. Further details are presented below.
3.1 Galaxy catalogs
@@
The redMaGiC galaxies are divided into two catalogs across five tomographic bins from $0.900.15<z<$ 0.90. Galaxies in the first three redshift bins ($0.650.15<z<$ 0.65) were drawn from a “high density” catalog, which is volume limited to a z-band luminosity of 0.5L*. The last two tomographic bins ($0.900.65<z<$ 0.90) are drawn from a “high luminosity high redshift” catalog, which is volume limited to a z-band luminosity of 1.0 L*. Because the galaxy population and selection function differ between the two samples, we perform our analysis on the high-density and high-redshift redMaGiC catalogs separately. An additional cut of $z>0.5$ is placed on the high-density redMaGiC sample to ensure adequate foreground-background separation. We use the $z<0.45$ population of the high-density redMaGiC catalog to investigate the extended CGM of luminous red galaxies.
@@
The redMaGiC catalogs do not contain photometry and must be joined with another catalog to obtain it. We use CosmoHub [^1] https://cosmohub.pic.es/home (J. Carretero et al., 2017; P. Tallada et al., 2020) to query the DES Y3 GOLD photometric catalog, using the galaxy selection criteria in I. Sevilla-Noarbe et al. (2021) to obtain a pure sample of galaxies with good photometry that can then be joined to the redMaGiC catalog. Additional selections to exclude galaxies with NaN, $-9999$, or $<28$ magnitudes in any bandpass are imposed; these cuts only eliminate an additional 1% of galaxies.
@@
The WISExSuperCOSMOS [^3] http://ssa.roe.ac.uk/WISExSCOS.html (or WISExSCOS) catalog combines mid-infrared data from the Wide-field Infrared Survey Explorer (WISE) and optical data from SuperCOSMOS scans of UKST/POSS-II photographic plates (M. Bilicki et al., 2016). The full catalog contains 20 million sources and covers the whole sky; our analysis is restricted to the subset of galaxies that overlap with the DES Y3 field and have corrected photometric redshifts $0.20<z<$ 0.2. An additional cut of $B^{\mathrm{cal}}_{\mathrm{corr}}>16$ is made to filter out stellar contaminants and very large galaxies that might be blended with background galaxies. Sky coverage of the WISExSCOS catalog, as well as redMaGiC, is shown in the top row of Figure 1.
Random (mock) galaxy catalogs, needed for the calculation of $\xi_{NA_{V}}(\theta)$, are provided for the two redMaGiC galaxy samples but not for WISExSCOS. Accordingly, we generate a random catalog for WISExSCOS by first sampling a large number of points uniformly in $(\theta,\sin\phi)$ then applying the WISExSCOS survey mask. The result is shown in the bottom left panel of Figure 1. Additional application of the DES Y3 survey mask produces the map shown in the bottom right panel of Figure 1.
@@
3.2 Star catalogs for systematics testing

### v2

@@ Abstract
One of the more surprising astrophysical discoveries of the last decade has been the presence of enormous quantities of dust at megaparsec distances from galaxies, which has important implications for galaxy evolution, the circumgalactic and intergalactic medium, and observational cosmology. In this work, we present a novel method for studying these vast halos of circumgalactic dust: a maximum-likelihood estimator for dust-induced extinction of background galaxies. This estimator can accommodate a broad range of archival photometric data and can incorporate different dust reddening prescriptions, making it applicable to diverse galaxy types and redshifts. We apply the estimator to the redMaGiC catalog of luminous red galaxies, selected for their tight dispersion in color and well-constrained photometric redshifts, and measure the resulting extinction as a function of projected distance from WISExSuperCOSMOS and redMaGiC foreground galaxies. We detect significant dust-induced extinction profiles extending to at least 1 megaparsec from galactic disks, with noticeable differences between star-forming and quiescent galaxies: star-forming galaxies exhibit a pronounced rise in extinction within the inner 50 kiloparsecs and a steep decline beyond 1 megaparsec, while the quiescent galaxies host little dust in the inner halo but have detectable extinction out to 30 megaparsecs. We test the robustness of our results using star catalogs and inverted foreground and background samples and find no evidence for significant systematic error. Our approach provides a powerful tool for studying the interplay between circumgalactic dust, galaxy evolution, and large-scale structure, with potential applications in a number of astrophysical subfields.
software: Astropy (Astropy Collaboration et al., 2013, 2018), NumPy (Harris et al., 2020), TreeCorr (Jarvis et al., 2004), healpy (Zonca et al., 2019)
@@ I Introduction
Most of the universe’s atoms are hidden from plain sight: galaxies account for less than 10% of baryons, while the remaining 90% reside in gaseous phases whose low densities make them difficult to observe. These include the circumgalactic medium (CGM), which permeates the dark matter halos hosting galaxies, and the more extended intergalactic medium (IGM) that connects them (Fukugita and Peebles, 2004; Shull et al., 2012). Combined, they are the reservoir for more than 85% of the baryons in the universe (Walter et al., 2020). Despite being the focus of intense theoretical and observational research, many open questions remain in the study of CGM and IGM, particularly regarding the origin, distribution, and survival of cool, condensed matter in these environments.
@@
One pressing issue is the existence of huge amounts of dust in the CGM and IGM—comparable to the amount of dust contained within galaxies themselves—which challenges models of the cosmic baryon cycle for several reasons (Veilleux et al., 2020; Péroux and Howk, 2020; Chen and Zahedy, 2024). First, dust is predominantly deposited into the interstellar medium (ISM) through supernovae and asymptotic giant branch star winds; second, the CGM and IGM are filled with mega-Kelvin gas, which should rapidly destroy dust grains via sputtering (Draine and Salpeter, 1979). While a number of mechanisms for dust deposition and replenishment have been proposed, including stellar radiation-pressure driven outflows, tidal stripping of interacting galaxies, supernova-driven outflows, and cool outflows, the exact coupling between the multi-phase CGM, IGM, and ISM remains unclear (Chen and Zahedy, 2024).
@@
Understanding this connection matters because the abundance and spatial extent of dust in the CGM is a key component of galaxy formation and evolution models. Dust also fundamentally limits the homogeneity of samples in wide-field surveys and introduces significant scatter to cosmological distance measurements, particularly supernovae. According to Rubin et al. (2023), dust is the second-largest source of uncertainty in Type Ia supernova cosmological parameter constraints, accounting for 11.5% of the total variance. Current analyses include intergalactic dust reddening in the systematic error budget, but because the magnitude of the effect is uncertain, they do not correct the distance measures themselves.
@@
While an early reddening-based indication of significant quantities of dust in galaxy halos (out to at least 60 kpc from galaxy disks) appeared in Zaritsky (1994), the existence of circumgalactic dust at megaparsec scales was first reported by Ménard, Scranton, Fukugita, and Richards (2010), hereafter M10. In that work, the authors reported the serendipitous discovery of a smooth extinction profile extending to 10 Mpc while measuring the magnification of 85,000 background quasars by 20 million galaxies in SDSS images. A similar measurement was carried out by Peek et al. (2015) using a background sample of luminous red galaxies (LRGs) as “standard crayons” (Peek and Graves, 2010), whose small color dispersion permitted detection of reddening by dust halos with a smaller sample size than M10. Since that time, the majority of investigations of the extended CGM have been spectroscopic, e.g., Zhu et al. (2014), Lan and Mo (2018), Zu (2021), and Chen et al. (2024). These studies typically use tracers like the equivalent width (EW) of Mg II absorption lines to probe the dynamics and shape of cool gas in the CGM of galaxies at $z>0.3$, where the Mg II $\lambda\lambda$ 2796, 2803 doublet is redshifted into the optical. In turn, this cool, metal-rich gas can be used to infer a dust abundance by relating the Mg II absorber equivalent width to a reddening value (Ménard and Fukugita, 2012). Far-infrared emission from CGM dust provides a more direct tracer of the dust content and thermal conditions, but current measurements are generally limited to $r\lesssim 100$ kpc from the optical radius of the galaxy (Zhang et al., 2016; Zu, 2021; Veilleux et al., 2025). Thus, with the notable exception of Ruoyi and Haibo (2020), who reported the detection of dust-induced reddening about M31 and M33 at projected distances greater than 100 kpc, almost no studies have replicated the original M10 circumgalactic dust measurement using extinction of background sources.
@@
These observational gaps highlight the need for a way to assess dust in the extended CGM without relying on atomic absorption lines unsuited for low-redshift galaxies, particularly given the evidence that the cosmic dust density changes significantly at $z<1$ (Péroux and Howk, 2020). Drawing inspiration from the “standard crayon” approach in Peek et al. (2015) and from analytical techniques ubiquitous in cosmology, we present a new technique to detect circumgalactic dust halos: a maximum-likelihood estimator for dust-induced extinction with a lower bound on the variance set by the Cramér-Rao bound. Application of the estimator is relatively simple, requiring only a photometric dataset and specification of a theoretical model for dust-induced reddening. Applying this estimator to the redMaGiC galaxy catalog, notable for its tight dispersion in color, and cross-correlating the inferred extinction with $z<0.2$ galaxies from the WISExSuperCOSMOS catalog, we find fresh evidence for circumgalactic dust halos at distances of 1 $h^{-1}\,$Mpc or greater from the galactic disk. We also analyze a sample of $0.15<z<0.45$ redMaGiC LRGs, and find distinct differences in the radial extinction profiles of star-forming and quiescent galaxies.
@@
The remainder of this paper is organized as follows. We describe our dust extinction estimator in Section II, provide details on datasets in Section III, and describe our methods in Section IV. Results are shown in Section V and discussed in Section VI. We conclude in Section VII.
II A maximum likelihood estimator for dust extinction
@@
A screen of dust absorbs and scatters incoming light, dimming and reddening the spectrum. The extinction $A_{\lambda}$ quantifies wavelength-dependent dimming of light due to dust; reddening is usually quantified by a color excess, most commonly the selective extinction of red versus blue light in the UBV photometric system: $E(\mathrm{B-V})$. The steepness of $A_{\lambda}$ as a function of wavelength is parametrized by the ratio of selective-to-total extinction in V-band: $R_{\mathrm{V}}=A_{\mathrm{V}}/E(\mathrm{B-V})$. For interstellar Milky Way dust, $R_{\mathrm{V}}\sim 3.1$; dense molecular clouds have $R_{\mathrm{V}}\sim 5$ (Fitzpatrick and Massa, 2007).
@@
Most photometric studies of the CGM trace dust using some variant of the selective extinction $E(\mathrm{B-V})$; here, we use maximum likelihood estimation (e.g., Barlow, 1993) to find a parameter that represents the best chance of finding dust-induced extinction based on a change in background galaxy colors. This estimator can be derived by maximizing the log-likelihood
@@
and the covariance of the data $\mathbf{C}$, calculated directly from the data vector $\mathbf{D}$, encapsulates both the magnitude measurement uncertainties and cosmic variance. The model $\mathbf{M}(\rho)$ describes the data $\mathbf{D}$ in terms of a parameter $\rho$ (here, $A_{\mathrm{V}}$); the model can be Taylor-expanded to obtain a simple linear approximation for the effect of $\rho$:
@@
In this expansion, $\mathrm{M_{0}}$ are the intrinsic colors of galaxies in the background galaxy sample, without any reddening from dust. The $\rho$ $d\mathrm{M}/d\rho$ term is the response of galaxies to a differential change in extinction; $d\mathrm{M}/d\rho$ is a fiducial extinction curve for dust with a given $R_{\mathrm{V}}$. With this assumption, the optimal estimator $\hat{\rho}$ for dust-induced extinction $\rho$ such that $\rho$ maximizes the likelihood in Equation 1 is given as:
@@
With the extinction estimator $\hat{\rho}$ in place, foreground galaxy number counts may be cross-correlated with the dust-induced extinction of background galaxies. Denoting $N(\theta_{1})$ as the galaxy count at a position $\theta_{1}$ on the sky and $A_{\mathrm{V}}(\theta_{2})$ as the extinction at position $\theta_{2}$, the cross-correlation may be expressed as the average deviations of $N$ and $A_{\mathrm{V}}$ from their mean values at two sky positions $\theta_{1}$ and $\theta_{2}$ separated by $\theta$:
$\xi_{NA_{\mathrm{V}}}(\theta)=\left<[N(\theta_{1})-\overline{N}]\,[A_{\mathrm{V}}(\theta_{2})-\overline{A_{\mathrm{V}}}]\right$> (6)
@@
The mean values, $\overline{N}$ and $\overline{A_{\mathrm{V}}}$, are the average foreground galaxy count and average extinction of background galaxies, respectively, over the entire survey area.
@@
As with any such analysis, the contribution of random correlations to $A_{\mathrm{V}}$ must be taken into account. We use the following estimator for the true galaxy-reddening cross-correlation $\xi_{NA_{\mathrm{V}}}(\theta)$:
$\xi_{NA_{\mathrm{V}}}(\theta)=D_{N}D_{A_{\mathrm{V}}}(\theta)-D_{N}R_{A_{\mathrm{V}}}(\theta)-R_{N}D_{A_{\mathrm{V}}}(\theta)+R_{N}R_{A_{\mathrm{V}}}(\theta)$ (7)
@@
$D_{N}D_{A_{\mathrm{V}}}$ is the signal from the correlation of the real foreground galaxy and background color catalogs; $D_{N}R_{A_{\mathrm{V}}}$ and $R_{N}D_{A_{\mathrm{V}}}$ are the correlations of real foreground with random background catalog and random foreground with real background catalogs, respectively; and $R_{N}R_{A_{\mathrm{V}}}$ is the correlation of the random foreground and background catalogs.
III Data
@@
We use the Dark Energy Survey (DES) Y3 redMaGiC catalogs (Rozo et al., 2016; Pandey et al., 2022) as background samples and the WISExSuperCOSMOS photometric redshift catalog (Bilicki et al., 2016), plus a low-redshift subset of redMaGiC galaxies, as foreground samples. In addition, star catalogs selected from the Gaia survey and DES Y3 are used to test for systematic error. Further details are presented below.
III.1 Galaxy catalogs
@@
The redMaGiC galaxies are divided into two catalogs across five tomographic bins from $0.15<z<0.90$. Galaxies in the first three redshift bins ($0.15<z<0.65$) were drawn from a “high density” catalog, which is volume limited to a z-band luminosity of 0.5L*. The last two tomographic bins ($0.65<z<0.90$) are drawn from a “high luminosity high redshift” catalog, which is volume limited to a z-band luminosity of 1.0 L*. Because the galaxy population and selection function differ between the two samples, we perform our analysis on the high-density and high-redshift redMaGiC catalogs separately. An additional cut of $z>0.5$ is placed on the high-density redMaGiC sample to ensure adequate foreground-background separation. We use the $z<0.45$ population of the high-density redMaGiC catalog to investigate the extended CGM of luminous red galaxies.
@@
The redMaGiC catalogs do not contain photometry and must be joined with another catalog to obtain it. We use CosmoHub [^1] https://cosmohub.pic.es/home (Carretero et al., 2017; Tallada et al., 2020) to query the DES Y3 GOLD photometric catalog, using the galaxy selection criteria in Sevilla-Noarbe et al. (2021) to obtain a pure sample of galaxies with good photometry that can then be joined to the redMaGiC catalog. Additional selections to exclude galaxies with NaN, 9999-9999, or <28<28 magnitudes in any bandpass are imposed; these cuts only eliminate an additional 1% of galaxies.
@@
The WISExSuperCOSMOS [^3] http://ssa.roe.ac.uk/WISExSCOS.html (or WISExSCOS) catalog combines mid-infrared data from the Wide-field Infrared Survey Explorer (WISE) and optical data from SuperCOSMOS scans of UKST/POSS-II photographic plates (Bilicki et al., 2016). The full catalog contains 20 million sources and covers the whole sky; our analysis is restricted to the subset of galaxies that overlap with the DES Y3 field and have corrected photometric redshifts $0<z<0.2$. An additional cut of $B^{\mathrm{cal}}_{\mathrm{corr}}>16$ is made to filter out stellar contaminants and very large galaxies that might be blended with background galaxies. Sky coverage of the WISExSCOS catalog, as well as redMaGiC, is shown in the top row of Figure 1.
Random (mock) galaxy catalogs, needed for the calculation of $\xi_{NA_{\mathrm{V}}}(\theta)$, are provided for the two redMaGiC galaxy samples but not for WISExSCOS. Accordingly, we generate a random catalog for WISExSCOS by first sampling a large number of points uniformly in $(\theta,\sin\phi)$ then applying the WISExSCOS survey mask. The result is shown in the bottom left panel of Figure 1. Additional application of the DES Y3 survey mask produces the map shown in the bottom right panel of Figure 1.
@@
III.2 Star catalogs for systematics testing